# Jigsaw-R1: A Study of Rule-based Visual Reinforcement Learning with Jigsaw Puzzles

### v2

@@ Abstract
The application of rule-based reinforcement learning (RL) to multimodal large language models (MLLMs) introduces unique challenges and potential deviations from findings in text-only domains, particularly for perception-heavy tasks. This paper provides a comprehensive study of rule-based visual RL, using jigsaw puzzles as a structured experimental framework. Jigsaw puzzles offer inherent ground truth, adjustable difficulty, and demand complex decision-making, making them ideal for this study. Our research reveals several key findings: Firstly, we find that MLLMs, initially performing near to random guessing on the simplest jigsaw puzzles, achieve near-perfect accuracy and generalize to complex, unseen configurations through fine-tuning. Secondly, training on jigsaw puzzles can induce generalization to other visual tasks, with effectiveness tied to specific task configurations. Thirdly, MLLMs can learn and generalize with or without explicit reasoning, though open-source models often favor direct answering. Consequently, even when trained for step-by-step reasoning, they can ignore the thinking process in deriving the final answer. Fourthly, we observe that complex reasoning patterns appear to be pre-existing rather than emergent, with their frequency increasing alongside training and task difficulty. Finally, our results demonstrate that RL exhibits more effective generalization than Supervised Fine-Tuning (SFT), and an initial SFT cold start phase can hinder subsequent RL optimization. Although these observations are based on jigsaw puzzles and may vary across other visual tasks, this research contributes a valuable piece of jigsaw to the larger puzzle of collective understanding rule-based visual RL and its potential in multimodal learning. The code is available at: https://github.com/zifuwanggg/Jigsaw-R1.
@@ 1 Introduction
Despite these advancements, the application of rule-based RL to multimodal contexts is still in its early stages. Unlike purely textual environments such as DeepSeek-R1, multimodal large language models (MLLMs) face the complex challenge of integrating and reasoning over both textual and visual information. This introduces unique difficulties and potential deviations from findings in purely linguistic domains (Gandhi et al., 2025; Guo et al., 2025; Hassid et al., 2025; Lee et al., 2025; Liu et al., 2025d; Marjanović et al., 2025; Xie et al., 2025b).
@@
This paper presents a comprehensive study of rule-based RL within the multimodal domain. Rather than relying on verifiable answers from existing MLLM benchmarks, we revisit a classic pretext task in computer vision: solving jigsaw puzzles (Carlucci et al., 2019; Chen et al., 2023; Doersch et al., 2015; Du et al., 2020; Kim et al., 2018; Noroozi & Favaro, 2016). This task (a visual illustration is presented Figure 1) offers a compelling testbed for studying rule-based visual RL for several reasons:
@@ 2.1 Jigsaw Puzzles
For instance, Doersch et al. (2015); Noroozi & Favaro (2016) propose pre-training Convolutional Neural Networks (CNNs) on jigsaw puzzles before fine-tuning them for downstream tasks such as image classification and object detection. Building on this, Kim et al. (2018) increase task complexity by integrating jigsaw puzzles with inpainting and colorization, demonstrating transferable representations for image classification and semantic segmentation. Moreover, Carlucci et al. (2019) employ jigsaw puzzles as a self-supervised regularization term, achieving strong domain generalization. Du et al. (2020) further advance this area by incorporating jigsaw puzzles into a progressive training pipeline for fine-grained classification. Most recently, Chen et al. (2023) investigate the efficacy of jigsaw puzzles in the context of vision transformers (ViTs), highlighting the importance for architecture-specific modifications. To the best of our knowledge, the adoption of jigsaw puzzles as pretext tasks in the era of LLMs remains an unexplored area.
@@ 2.2 Rule-based Reinforcement Learning in Textual Domains
Beyond the generalization capabilities, DeepSeek-R1 also demonstrates a natural achievement of test-time scaling (Snell et al., 2025), evidenced by increased completion lengths and the sudden emergence of complex reasoning patterns, a phenomenon termed the aha moment. However, Hassid et al. (2025); Xie et al. (2025b) find that longer responses do not guarantee better reasoning and Liu et al. (2025d) challenge the notion of sudden emergence, positing that these complex behaviors might be inherent in the base model rather than appearing abruptly.
Furthermore, its reasoning processes reportedly exhibit human-like language processing characteristics (Marjanović et al., 2025). Investigating these further, Gandhi et al. (2025) draw connections to human psychology, developing a framework that categorizes these reasoning patterns into four key cognitive behaviors. Their findings suggest that Qwen2.5 (Qwen et al., 2025) naturally exhibits these behaviors, whereas Llama3.2 (Grattafiori et al., 2024) initially does not. Adding another dimension to this discussion, Lee et al. (2025) observe that the data format has a significantly greater impact on reasoning behavior than the data domain. Their analysis reveals that models trained on multiple-choice data produce concise answers, whereas those trained on free-form data tend to be more verbose.
2.3 Rule-based Reinforcement Learning without Labeled Datasets
While effective, rule-based RL often requires extensive human efforts to curate verifiable data. To mitigate this, Xie et al. (2025b) leverage a logic puzzle called Knights and Knaves (K&K), which allows for controllable difficulty and algorithmically generated ground truth. They demonstrate that models trained on this synthetic dataset can develop complex reasoning skills transferrable to mathematical tasks. Subsequently, Chen et al. (2025b); Liu et al. (2025a) extend this approach to a broader range of textual puzzles, including Sudoku and game of 24. For a greater level of autonomy, Zhao et al. (2025a) introduce a framework where a single model learns to propose tasks that maximize its own learning progress and improves reasoning by solving them, without relying on any external data.
@@ 2.4 Rule-based Reinforcement Learning in Multimodal Domains
A significant line of exploration, motivated by the success of DeepSeek-R1 in mathematical reasoning, involves adapting these rule-based RL techniques to multimodal mathematical tasks (Chen et al., 2025a; Deng et al., 2025; Liu et al., 2025b; Meng et al., 2025; Peng et al., 2025; Wang et al., 2025b; Yang et al., 2025). These efforts have consistently demonstrated strong out-of-domain generalization. Notably, studies by Chen et al. (2025a); Meng et al. (2025) have documented the emergence of the aha moment—the spontaneous exhibition of sophisticated reasoning patterns like self-correction—arising from end-to-end RL training.
However, the landscape appears different for visual perception tasks, including visual classification, visual grounding and spatial reasoning. While research in this area (Bai et al., 2025b; Chen et al., 2025d; Lai et al., 2025; Li et al., 2025b; Liao et al., 2025; Liu et al., 2025e; c; Shen et al., 2025; Yu et al., 2025a; Zhou et al., 2025) also show robust out-of-distribution generalization, the aha moment has not been observed when employing instruction-tuned models. Indeed, the necessity of explicit reasoning steps for these perception-intensive tasks has been questioned, as direct answers may often suffice (Jiang et al., 2025). In fact, work by Lai et al. (2025); Li et al. (2025b); Yu et al. (2025a) suggest that training models for direct answering often leads to superior performance compared to models trained for explicit step-by-step reasoning.
Figure 1: An illustration of the task design.
@@ 3 Task Design
This section outlines the formulation of jigsaw puzzles in a format suitable for processing by MLLMs. Subsequently, we introduce a rule-based reward system designed for RL training. All specific prompts are detailed in Appendix E, and Figure 1 provides a conceptual overview.
3.1 Jigsaw Puzzle Images
The creation of jigsaw puzzles begins with an input image. The image is first partitioned into an $m{\scriptstyle\times}n$ grid of patches and the task’s difficulty can be readily adjusted by varying the values of $m$ and $n$. Optionally, a masked region can be added between patches to highlight the grid layout. If the image’s height is not perfectly divisible by $m$, or its width by $n$, the image is trimmed from the bottom or right edges to ensure its dimensions are exact multiples of the patch count. Subsequently, these patches are randomly shuffled to create the jigsaw puzzle images. To uniquely identify each patch’s location within this grid, position indices are assigned sequentially in row-major order, from 1 (top-left) to $mn$ (bottom-right).
@@ 3.2 Question Types
Based on these shuffled images, we formulate distinct question types. These questions either directly assess the MLLM’s ability to reconstruct the original image or require it to reason about the relative positions between the shuffled patches in the initial image. An additional question type, assessing both spatial reasoning and visual grounding, is presented in Appendix B.
@@ 3.4 Rule-based Rewards
The format reward is 0.5 if the output adheres to all these requirements, and 0 otherwise.
@@ 4.3 Implementation Details
We use GRPO (Shao et al., 2024) as the reinforcement learning algorithm. The GRPO iteration $\mu=1$, the KL efficient $\beta=0.04$ and the clipping value $\epsilon=0.2$. The number of training steps is 1,000 for thinking and 2,000 for non-thinking. In each training step, 64 unique prompts are processed, with each prompt being sampled 8 times to calculate the advantages. The sampling temperature is set to 1, and top-k sampling is used with $k=50$. The learning rate initiates at 1e-6 and linearly decays to 0.
@@ 5 Experiments
This section presents results designed to address the proposed research questions. The main paper focuses on instruction-tuned models; for a discussion regarding Qwen2-VL-2B-Base, please refer to Appendix C.
@@
To answer the question, we first train models on 2x1 jigsaw puzzles using the training split of the COCO dataset. To introduce task diversity for non-square puzzles, the piece order is randomly shuffled in 50% of instances (e.g. 2x1 becomes 1x2 and vice verse). We then evaluate model performance on the same question type but with varying puzzle sizes, utilizing the test split of the COCO dataset. Evaluation results for pair questions are shown in Table 1, and comprehensive results for full questions are provided in Appendix A. The training dynamics for pair jigsaw puzzles are illustrated in Figure 2.
Finding 1.1: MLLMs struggle with jigsaw puzzles before fine-tuning. As demonstrated in Tables 1 and 7, jigsaw puzzles are notably difficult for MLLMs without task-specific training. Prior to fine-tuning, even powerful proprietary models perform at levels comparable to random guessing, struggling even with the simplest jigsaw puzzles (i.e., 2x1).
Finding 1.2: MLLMs exhibit efficient learning and generalization for jigsaw puzzles after fine-tuning. Despite the initial difficulty, MLLMs show a strong capacity to learn and solve these puzzles after fine-tuning. For example, the reward progression for Qwen2.5-VL-3B, depicted in Figure 2, indicates rapid convergence to near-perfect accuracy. Notably, models trained exclusively on 2x1 jigsaw puzzles successfully generalize their learned abilities to larger puzzle sizes beyond the training distribution (e.g., 3x1).
@@
To address this question, we evaluate the performance of models trained on jigsaw puzzles across several downstream tasks, including CV-Bench, MMVP, SAT, and Super-CLEVR. Further details regarding these datasets can be found in Section 4.1.
Finding 2.1: Jigsaw puzzles generalize. Our primary investigation (Table 2) reveals that models trained on jigsaw puzzles generally achieve improved performance on downstream tasks, indicating robust generalization. Notably, despite being trained exclusively on the COCO dataset for jigsaw puzzles, these models successfully adapt to spatial reasoning tasks on synthetic image datasets such as SAT and Super-CLEVR. Nevertheless, the improvements seen in thinking models might be superficial. They learn to neglect the reasoning process (further discussed in Research Question #3), and their performance after fine-tuning remains inferior to that of non-thinking models that have not undergone fine-tuning.
@@
Puzzle size: The choice of jigsaw puzzle size significantly affects downstream performance (Table 3). For the non-thinking setting, training on a larger, more challenging jigsaw puzzle size leads to better generalization. Furthermore, employing a curriculum learning approach that mixes different puzzle sizes (e.g. 3x1$\rightarrow$4x1) proves more effective than training exclusively with a single size.
Question type: Pair jigsaw puzzles result in superior generalization on downstream tasks compared to full jigsaw puzzles (Table 4). We believe that this advantage stems from the analogy of pair jigsaw puzzles to downstream tasks (e.g. requiring models to answer multi-choice questions and explicitly asking them to reason about spatial relationships between visual elements).
Training dataset: As demonstrated in Table 5, aligning the training dataset with the target domain yields improved performance. For example, training directly on the SAT dataset enhances performance on SAT tasks. Given that jigsaw puzzles are label-free, it is even feasible to train on the test set of SAT for further performance gains.
@@
Table 3: Averaged downstream task performance of Qwen2.5-VL-3B when trained on different jigsaw puzzle sizes. 2x1$\rightarrow$3x1 (3x1$\rightarrow$4x1): Mixing 2x1 and 3x1 (3x1 and 4x1) jigsaw puzzles in a curriculum setting. For thinking and non-thinking of the same model, the better result is underlined.
@@
2x1 3x1 4x1 2x2 3x1$\rightarrow$4x1
$\underline{73.18}$ $\underline{74.95}$ $\underline{73.03}$ $\underline{72.68}$ $\underline{\textbf{75.29}}$
@@
Finding 3.1: MLLMs can learn from jigsaw puzzles and generalize to downstream tasks with or without explicit reasoning. As demonstrated by the jigsaw puzzle results (Tables 1 and 7), models can effectively learn with rule-based visual RL, whether or not an explicit reasoning chain is generated. Importantly, these learned capabilities can be generalized to various downstream tasks (Table 2).
Finding 3.2: Open-source MLLMs often benefit from direct answering, while proprietary models tend to perform better with explicit reasoning. Consistent with observations in (Li et al., 2025b; Jiang et al., 2025), our results on jigsaw puzzles (Tables 1 and 7) and downstream tasks (Table 2) confirm that open-source models tend to achieve stronger results when prompted to output the answer directly. Indeed, even after fine-tuning, models that adopt a reasoning process show weaker generalization on downstream tasks than direct-answering models that have not undergone fine-tuning. Conversely, we find that proprietary models generally demonstrate improved performance when an explicit reasoning process precedes the final answer. It is important to note, however, that this does not necessarily mean proprietary models are inherently stronger. For example, Claude 3.5 Haiku performs comparably to Qwen2.5-VL-3B on downstream tasks.
Finding 3.3: MLLMs can neglect the reasoning process after fine-tuning. As illustrated in Figure 2 (Right), the completion length of InternVL2.5-2B significantly decreases during training. This occurs because the model increasingly circumvents step-by-step reasoning and often includes only the final answer in its thinking process (see examples in Appendix G). Conversely, while Qwen models do present explicit reasoning steps, these steps may not consistently inform the derivation of the final answer (examples are provided in Appendix F). To quantify this, we utilized GPT-4.1 to assess the consistency between the reasoning process of Qwen2.5-VL-3B and its final answer. As Figure 3 (Left) demonstrates, although the model’s final answer becomes more accurate with training, its reasoning chain becomes progressively more inconsistent.
Takeaways #3. MLLMs can learn and generalize, irrespective of whether an explicit reasoning process is included. Nevertheless, open-source MLLMs usually excel at direct answering. Consequently, even when trained to utilize step-by-step reasoning, they may ignore the thinking process when deiriving the final answer.
@@
Finding 4.1: Complex reasoning patterns are pre-existing in MLLMs. The aha moment is often associated with the emergence of complex reasoning patterns Guo et al. (2025). While contemporary studies focusing on perception-heavy tasks like visual classification, visual grounding and spatial reasoning (Bai et al., 2025b; Chen et al., 2025d; Lai et al., 2025; Li et al., 2025b; Liao et al., 2025; Liu et al., 2025e; c; Shen et al., 2025; Yu et al., 2025a; Zhou et al., 2025) typically do not observe these patterns in instruction-tuned models, our investigation into jigsaw puzzles reveals a distinct phenomenon. Although the completion length is not increasing, as illustrated in Figure 2 (Right), we find that all these models, including InternVL2.5-2B and also Qwen2-VL-2B-Base (see Appendix C), exhibit complex reasoning patterns, such as verification and backtracking, even before training starts. Indeed, throughout the training process, we successfully identify all four cognitive behaviors as defined in (Gandhi et al., 2025).
@@
Finding 4.2: Complex reasoning patterns evolve during fine-tuning. To monitor the evolution of these behaviors, we track the frequency of keywords indicative of backtracking and backward chaining (detailed in Appendix D). As depicted in Figure 3 (Middle), the occurrence of these keywords demonstrates a steady and significant increase throughout the training process.
Finding 4.3: Complex reasoning patterns emerge more frequently with harder jigsaw puzzles. To further investigate these reasoning patterns, we plot the frequency of these keywords as the Qwen2.5-VL-3B model is trained on jigsaw puzzles of varying sizes, as shown in Figure 3 (Right). Our analysis reveals a clear trend: the frequency of these keywords increases when the model is trained on more challenging (i.e., larger) jigsaw puzzles.
@@
This section evaluates the generalization capabilities of SFT in comparison to RL. For these experiments, SFT data for thinking is curated via rejection sampling, while non-thinking utilizes ground-truth data.
Finding 5.1: SFT exhibits weaker generalization than RL. As demonstrated in Table 6, applying SFT to either the reasoning chain (thinking) or directly to ground-truth answers (non-thinking) can yield some generalization. However, it is generally less effective than RL.
Finding 5.2: A cold start phase with SFT preceding RL can be detrimental. Compared to a single-stage RL process, a two-stage pipeline that incorporates a cold start phase with SFT prior to RL can help models learn specific output formats (Guo et al., 2025). However, this is not essential in our experimental setting, as indicated by the rapid increase in rewards (Figure 2). More importantly, we observe that this cold start phase can diminish the effectiveness of subsequent RL optimization (Table 6).

### v3

@@ Abstract
The application of rule-based reinforcement learning (RL) to multimodal large language models (MLLMs) introduces unique challenges and potential deviations from findings in text-only domains, particularly for perception-heavy tasks. This paper provides a comprehensive study of rule-based visual RL, using jigsaw puzzles [^2] For a preliminary study with another pretext task, image rotation, please refer to Appendix C. as a structured experimental framework. Jigsaw puzzles offer inherent ground truth, adjustable difficulty, and demand complex decision-making, making them ideal for this study. Our research reveals several key findings: Firstly, we find that MLLMs, initially performing near to random guessing on the simplest jigsaw puzzles, achieve near-perfect accuracy and generalize to complex, unseen configurations through fine-tuning. Secondly, training on jigsaw puzzles can induce generalization to other visual tasks, with effectiveness tied to specific task configurations. Thirdly, MLLMs can learn and generalize with or without explicit reasoning, though open-source models often favor direct answering. Consequently, even when trained for step-by-step reasoning, they can ignore the thinking process in deriving the final answer. Fourthly, we observe that complex reasoning patterns appear to be pre-existing rather than emergent, with their frequency increasing alongside training and task difficulty. Finally, our results demonstrate that RL exhibits more effective generalization than Supervised Fine-Tuning (SFT), and an initial SFT cold start phase can hinder subsequent RL optimization. Although these observations are based on jigsaw puzzles and may vary across other visual tasks, this research contributes a valuable piece of jigsaw to the larger puzzle of collective understanding rule-based visual RL and its potential in multimodal learning. The code is available at: https://github.com/zifuwanggg/Jigsaw-R1.
Zifu Wang∗1   Junyi Zhu∗1   Bo Tang∗2,3   Zhiyu Li3   Feiyu Xiong3
Jiaqian Yu4   Matthew B. Blaschko1
1ESAT-PSI, KU Leuven
2University of Science and Technology of China
3Memory Tensor
4Samsung R&D Institute China
11footnotetext: Equal contribution. Correspondence to: zifu.wang@kuleuven.be
@@ 1 Introduction
Despite these advancements, the application of rule-based RL to multimodal contexts is still in its early stages. Unlike purely textual environments such as DeepSeek-R1, multimodal large language models (MLLMs) face the complex challenge of integrating and reasoning over both textual and visual information. This introduces unique difficulties and potential deviations from findings in purely linguistic domains (Gandhi et al., 2025; Guo et al., 2025; Hassid et al., 2025; Lee et al., 2025; Liu et al., 2025e; Marjanović et al., 2025; Xie et al., 2025b).
@@
This paper presents a comprehensive study of rule-based RL within the multimodal domain. Rather than relying on verifiable answers from existing MLLM benchmarks, we revisit a classic pretext task in computer vision: solving jigsaw puzzles (Carlucci et al., 2019; Chen et al., 2023; Doersch et al., 2015; Du et al., 2020; Kim et al., 2018; Noroozi & Favaro, 2016). This task (a visual illustration is presented Figure˜1) offers a compelling testbed for studying rule-based visual RL for several reasons:
@@ 2.1 Jigsaw Puzzles
For instance, Doersch et al. (2015); Noroozi & Favaro (2016) propose pre-training Convolutional Neural Networks (CNNs) on jigsaw puzzles before fine-tuning them for downstream tasks such as image classification and object detection. Building on this, Kim et al. (2018) increase task complexity by integrating jigsaw puzzles with inpainting and colorization, demonstrating transferable representations for image classification and semantic segmentation. Moreover, Carlucci et al. (2019) employ jigsaw puzzles as a self-supervised regularization term, achieving strong domain generalization. Du et al. (2020) further advance this area by incorporating jigsaw puzzles into a progressive training pipeline for fine-grained classification. Most recently, Chen et al. (2023) investigate the efficacy of jigsaw puzzles in the context of vision transformers (ViTs), highlighting the importance for architecture-specific modifications. Nevertheless, to the best of our knowledge, the adoption of jigsaw puzzles as pretext tasks in the era of LLMs remains an unexplored area.
@@ 2.2 Rule-based Reinforcement Learning in Textual Domains
Beyond the generalization capabilities, the reasoning chains produced by models trained with rule-based RL have been the subject of extensive study. For example, DeepSeek-R1 demonstrates a natural achievement of test-time scaling (Snell et al., 2025), evidenced by increased completion lengths and the sudden emergence of complex reasoning patterns, a phenomenon termed the aha moment. However, these findings are debated. Hassid et al. (2025); Xie et al. (2025b) contend that longer responses do not guarantee better reasoning and Liu et al. (2025e) challenge the notion of sudden emergence, positing that these complex behaviors might be inherent in the base model rather than appearing abruptly. Separately, the reasoning processes within DeepSeek-R1 have also been noted for exhibiting human-like language processing characteristics (Marjanović et al., 2025). Investigating these further, Gandhi et al. (2025) draw connections to human psychology, developing a framework that categorizes these reasoning patterns into four key cognitive behaviors. Their findings suggest that Qwen2.5 (Qwen et al., 2025) naturally exhibits these behaviors, whereas Llama3.2 (Grattafiori et al., 2024) initially does not.
@@ 2.3 Rule-based Reinforcement Learning in Multimodal Domains
A significant line of exploration, motivated by the success of DeepSeek-R1 in mathematical reasoning, involves adapting rule-based RL techniques to multimodal mathematical tasks (Chen et al., 2025a; Deng et al., 2025; Liu et al., 2025c; Meng et al., 2025; Peng et al., 2025; Wang et al., 2025b; Yang et al., 2025). These efforts have consistently demonstrated strong out-of-domain generalization. Notably, studies by Chen et al. (2025a); Meng et al. (2025) have documented the emergence of the aha moment—the spontaneous exhibition of sophisticated reasoning patterns like self-correction—arising from end-to-end RL training.
However, the landscape appears different for visual perception tasks, including visual classification, visual grounding and spatial reasoning. While research in this area (Bai et al., 2025b; Chen et al., 2025d; Lai et al., 2025; Li et al., 2025c; Liao et al., 2025; Liu et al., 2025f; d; Shen et al., 2025; Yu et al., 2025a; Zhou et al., 2025) also show robust out-of-distribution generalization, the aha moment has not been observed when employing instruction-tuned models. Indeed, the necessity of explicit reasoning steps for these perception-intensive tasks has been questioned, as direct answers may often suffice (Jiang et al., 2025). In fact, work by Lai et al. (2025); Li et al. (2025c); Yu et al. (2025a) suggest that training models for direct answering often leads to superior performance compared to models trained for explicit step-by-step reasoning.
2.4 Rule-based Reinforcement Learning without Labeled Datasets
While effective, rule-based RL often requires extensive human efforts to curate verifiable data. To mitigate this, Xie et al. (2025b) leverage a logic puzzle called Knights and Knaves (K&K), which allows for controllable difficulty and algorithmically generated ground truth. They demonstrate that models trained on this synthetic dataset can develop complex reasoning skills transferrable to mathematical tasks. Building on this, Chen et al. (2025b); Liu et al. (2025a); Stojanovski et al. (2025) extend it to a wider array of textual puzzles, such as Sudoku and the game of 24. In a similar spirit, Feng et al. (2025); Tong et al. (2025) design several synthetic visual puzzles, demonstrating generalized reasoning abilities in MLLMs. For a greater level of autonomy, Zhao et al. (2025a) introduce a framework where a single model learns to propose tasks that maximize its own learning progress and improves reasoning by solving them, without relying on any external data.
Figure 1: An overview of the task design. An original image is divided into an $m{\scriptstyle\times}n$ grid of patches, which are then randomly shuffled to create the jigsaw image. We consider two question types (full or pair) and two prompting strategies (thinking or non-thinking).
@@ 3 Task Design
This section outlines the formulation of jigsaw puzzles in a format suitable for processing by MLLMs. Subsequently, we introduce a rule-based reward system designed for RL training. All specific prompts are detailed in Appendix˜I, and Figure˜1 provides a conceptual overview.
3.1 Jigsaw Images
The creation of jigsaw puzzles begins with an input image. The image is first partitioned into an $m{\scriptstyle\times}n$ grid of patches and the task’s difficulty can be readily adjusted by varying the values of $m$ and $n$. Optionally, a masked region can be added between patches to highlight the grid layout. If the image’s height is not perfectly divisible by $m$, or its width by $n$, the image is trimmed from the bottom or right edges to ensure its dimensions are exact multiples of the patch count. Subsequently, these patches are randomly shuffled to create the jigsaw images. To uniquely identify each patch’s location within this grid, position indices are assigned sequentially in row-major order, from 1 (top-left) to $mn$ (bottom-right).
@@ 3.2 Question Types
Based on these shuffled images, we formulate distinct question types. These questions either directly assess the MLLM’s ability to reconstruct the original image or require it to reason about the relative positions between the shuffled patches in the initial image. An additional question type, assessing both spatial reasoning and visual grounding, is presented in Appendix˜A.
@@ 3.4 Rule-based Rewards
The format reward is 0.5 if the output adheres to all these requirements, and 0 otherwise. In Appendix˜B, we present an ablation study of our reward design, comparing binary versus fractional accuracy rewards for full questions and analyzing the impact of format reward weights.
@@ 4.3 Implementation Details
We use GRPO (Shao et al., 2024) as the reinforcement learning algorithm. The GRPO iteration $\mu=1$, the KL efficient $\beta=0.04$ and the clipping value $\epsilon=0.2$. Given that thinking is substantially more computationally expensive, we perform 1,000 training steps for it, compared to 2,000 steps for non-thinking. More details on training costs are in Appendix˜D, and an ablation on the number of training steps is in Appendix˜E. In each training step, 64 unique prompts are processed, with each prompt being sampled 8 times to calculate the advantages. The sampling temperature is set to 1, and top-k sampling is used with $k=50$. The learning rate initiates at 1e-6 and linearly decays to 0.
As for SFT, we prepare the thinking data using the same prompt (as provided in Appendix˜I) to instruct the model to include a reasoning process. We then apply rejection sampling to retain only those instances where the thinking process correctly leads to the final answer. This complete output, including both the reasoning chain and the final answer, is subsequently used for fine-tuning. For non-thinking, we fine-tune the model directly on ground-truth answers. Both configurations are trained for 1,000 steps, with a batch size of 512. All other hyperparameters such as the learning rate are identical to those used in RL.
@@ 5 Experiments
This section presents results designed to address the proposed research questions. The main paper focuses on instruction-tuned models; for a discussion regarding Qwen2-VL-2B-Base, please refer to Appendix˜F.
@@
To answer the question, we first train models on 2x1 jigsaw puzzles using the training split of the COCO dataset. To introduce task diversity for non-square puzzles, the piece order is randomly shuffled in 50% of instances (e.g. 2x1 becomes 1x2 and vice verse). We then evaluate model performance on the same question type but with varying puzzle sizes, utilizing the test split of the COCO dataset. Evaluation results for pair questions are shown in Table˜1, and the corresponding training dynamics are illustrated in Figure˜2. Additional results, including full jigsaw puzzles and few-shot learning, are provided in Appendix˜A and Appendix˜G, respectively.
Finding 1.1: MLLMs struggle with jigsaw puzzles before fine-tuning. As demonstrated in Tables˜1 and 7, jigsaw puzzles are notably difficult for MLLMs without task-specific training. Prior to fine-tuning, even powerful proprietary models perform at levels comparable to random guessing, struggling even with the simplest jigsaw puzzles (i.e., 2x1).
Finding 1.2: MLLMs exhibit efficient learning and generalization for jigsaw puzzles after fine-tuning. Despite the initial difficulty, MLLMs show a strong capacity to learn and solve these puzzles after fine-tuning. For example, the reward progression for Qwen2.5-VL-3B, depicted in Figure˜2, indicates rapid convergence to near-perfect accuracy. Notably, models trained exclusively on 2x1 jigsaw puzzles successfully generalize their learned abilities to larger puzzle sizes beyond the training distribution (e.g., 3x1).
@@
To address this question, we evaluate the performance of models trained on jigsaw puzzles across several downstream tasks, including CV-Bench, MMVP, SAT, and Super-CLEVR. Further details regarding these datasets can be found in Section˜4.1.
Finding 2.1: Jigsaw puzzles generalize. Our primary investigation (Table˜2) reveals that models trained on jigsaw puzzles generally achieve improved performance on downstream tasks, indicating robust generalization. Notably, despite being trained exclusively on the COCO dataset for jigsaw puzzles, these models successfully adapt to spatial reasoning tasks on synthetic image datasets such as SAT and Super-CLEVR. Nevertheless, the improvements seen in thinking models might be superficial. They learn to neglect the reasoning process (further discussed in Research Question #3), and their performance after fine-tuning remains inferior to that of non-thinking models that have not undergone fine-tuning.
@@
[…]
Question type: Pair jigsaw puzzles result in superior generalization on downstream tasks compared to full jigsaw puzzles (Table˜4). We believe that this advantage stems from the analogy of pair jigsaw puzzles to downstream tasks (e.g. requiring models to answer multi-choice questions and explicitly asking them to reason about spatial relationships between visual elements).
Training dataset: As demonstrated in Table˜5, aligning the training dataset with the target domain yields improved performance. For example, training directly on the SAT dataset enhances performance on SAT tasks. Given that jigsaw puzzles are label-free, it is even feasible to train on the test set of SAT for further performance gains.
@@
[…]
@@
[…]
$\underline{69.59}$ $\underline{73.18}$ $\underline{74.95}$ $\underline{73.03}$ $\underline{72.68}$ $\underline{\textbf{75.29}}$
@@
In this section, we compare models that employ a reasoning chain against those that provide an answer directly.
Finding 3.1: MLLMs can learn from jigsaw puzzles and generalize to downstream tasks with or without explicit reasoning. As demonstrated by the jigsaw puzzle results (Tables˜1 and 7), models can effectively learn with rule-based visual RL, whether or not an explicit reasoning chain is generated. Importantly, these learned capabilities can be generalized to various downstream tasks (Table˜2).
Finding 3.2: Open-source MLLMs often benefit from direct answering, while proprietary models tend to perform better with explicit reasoning. Consistent with observations in (Li et al., 2025c; Jiang et al., 2025), our results on jigsaw puzzles (Tables˜1 and 7) and downstream tasks (Table˜2) confirm that open-source models tend to achieve stronger results when prompted to output the answer directly. Indeed, even after fine-tuning, models that adopt a reasoning process show weaker generalization on downstream tasks than direct-answering models that have not undergone fine-tuning. Conversely, we find that proprietary models generally demonstrate improved performance when an explicit reasoning process precedes the final answer, which mirrors findings on multimodal reasoning tasks (Hao et al., 2025). It is important to note, however, that this does not necessarily mean proprietary models are inherently stronger. For example, Claude 3.5 Haiku performs comparably to Qwen2.5-VL-3B on downstream tasks.
Finding 3.3: MLLMs can neglect the reasoning process after fine-tuning. As illustrated in Figure˜2 (Right), the completion length of InternVL2.5-2B significantly decreases during training. This occurs because the model increasingly circumvents step-by-step reasoning and often includes only the final answer in its thinking process (see examples in Appendix˜J). Conversely, while Qwen models do present explicit reasoning steps, these steps may not consistently inform the derivation of the final answer (examples are provided in Appendix˜J). To quantify this, we utilized GPT-4.1 to assess the consistency between the reasoning process of Qwen2.5-VL-3B and its final answer. As Figure˜3 (Left) demonstrates, although the model’s final answer becomes more accurate with training, its reasoning chain becomes progressively more inconsistent.
Takeaways #3. MLLMs can learn and generalize, irrespective of whether an explicit reasoning process is included. Nevertheless, open-source MLLMs usually excel at direct answering. Consequently, even when trained to include step-by-step reasoning, they may ignore the thinking process when deriving the final answer.
@@
This section focuses exclusively on the reasoning chains produced by thinking models. By design, non-thinking models provide a direct answer without generating an explicit reasoning process, so they are not included in this part of the analysis.
Finding 4.1: Complex reasoning patterns are pre-existing in MLLMs. The aha moment is often associated with the emergence of complex reasoning patterns (Guo et al., 2025). While contemporary studies focusing on perception-heavy tasks like visual classification, visual grounding and spatial reasoning (Bai et al., 2025b; Chen et al., 2025d; Lai et al., 2025; Li et al., 2025c; Liao et al., 2025; Liu et al., 2025f; d; Shen et al., 2025; Yu et al., 2025a; Zhou et al., 2025) typically do not observe these patterns in instruction-tuned models, our investigation into jigsaw puzzles reveals a distinct phenomenon. Although the completion length is not increasing, as illustrated in Figure˜2 (Right), we find that all these models, including InternVL2.5-2B and also Qwen2-VL-2B-Base (see Appendix˜F), exhibit complex reasoning patterns, such as verification and backtracking, even before training starts. Indeed, throughout the training process, we successfully identify all four cognitive behaviors as defined in (Gandhi et al., 2025).
@@
Finding 4.2: Complex reasoning patterns evolve during fine-tuning. To monitor the evolution of these behaviors, we track the frequency of keywords indicative of backtracking and backward chaining (detailed in Appendix˜H). As depicted in Figure˜3 (Middle), the occurrence of these keywords demonstrates a steady and significant increase throughout the training process. In addition to keyword matching, we also utilize GPT-4.1 to identify these behaviors in the reasoning chains. Further details are provided in Appendix˜H.
Finding 4.3: Complex reasoning patterns emerge more frequently with harder jigsaw puzzles. To further investigate these reasoning patterns, we plot the frequency of these keywords as the Qwen2.5-VL-3B model is trained on jigsaw puzzles of varying sizes, as shown in Figure˜3 (Right). Our analysis reveals a clear trend: the frequency of these keywords increases when the model is trained on more challenging (i.e., larger) jigsaw puzzles.
@@
This section evaluates the generalization capabilities of SFT in comparison to RL. For these experiments, SFT data for thinking is curated via rejection sampling, while non-thinking utilizes ground-truth data. Please refer to Section˜4.1 for more details.
Finding 5.1: SFT exhibits weaker generalization than RL. As demonstrated in Table˜6, applying SFT to either the reasoning chain (thinking) or directly to ground-truth answers (non-thinking) can yield some generalization. However, it is generally less effective than RL.
Finding 5.2: A cold start phase with SFT preceding RL can be detrimental. Compared to a single-stage RL process, a two-stage pipeline that incorporates a cold start phase with SFT prior to RL can help models learn specific output formats (Guo et al., 2025). However, this is not essential in our experimental setting, as indicated by the rapid increase in rewards (Figure˜2). More importantly, we observe that this cold start phase can diminish the effectiveness of subsequent RL optimization (Table˜6).